# Are penalty shootouts better than a coin toss? Evidence from international club football in Europe

### v3

@@ Abstract
Penalty shootouts play an important role in the knockout stage of major football tournaments, especially since the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule in its club competitions. Inspired by this rule change, our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, we find no evidence for the effect of the kicking order, the field of the match, or psychological momentum. In contrast to previous results, stronger teams, defined first by Elo ratings, do not perform better than their weaker opponents. Consequently, penalty shootouts seem to be close to a coin toss in top European club football.
@@ 1 Introduction
Penalty shootouts are the subject of serious academic interest since Apesteguia and Palacios-Huerta, (2010) has appeared in a leading economic journal, the American Economic Review, one and a half decades ago (see Section 2). In particular, the authors found a systematic first-mover advantage; the team kicking the first penalty has a significantly higher probability of winning the penalty shootout than its opponent.
The present paper contributes to this discussion by investigating four research questions:
  1. 1.
    Does the shooting order affect the outcome of a penalty shootout?
  2. 2.
    Does the field of the match affect the outcome of a penalty shootout?
  3. 3.
    Does psychological momentum affect the outcome of a penalty shootout?
  4. 4.
    Does the strength of the teams affect the outcome of a penalty shootout?
@@
All of these issues have partially been investigated in the existing literature, as revealed by Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, psychological momentum enjoyed by the comeback team is identified in a novel way, according to which team scored the last goal. Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used either the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), or betting odds (Wunderlich et al.,, 2020; Pipke,, 2025) for this purpose.
@@
As we show, the outcomes of the 268 penalty shootouts played in UEFA club competitions between 2000 and 2025 seem to be entirely unpredictable. This contradicts the results of some previous works, which uncover a first-mover advantage (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020) and the positive impact of psychological momentum (Krumer,, 2021). Furthermore, according to our knowledge, all existing studies agree that stronger teams are more likely to win in a shootout (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025), which cannot be observed in UEFA club competitions.
@@
Our findings could be especially important for the Union of European Football Associations (UEFA) when they decide on rule changes in the future. In 2021, UEFA abolished the so-called away goals rule, which awarded the prize (qualification for the next round) in a two-legged match to the team that scored more away goals. This change has increased the probability of reaching a penalty shootout by approximately ten percentage points, from below 5% to nearly 15% (Forrest et al.,, 2025). In addition, UEFA has recently given serious consideration to removing extra time in order to ease the burden of the best players (Ames,, 2025). Since the outcome of a penalty shootout approximates a coin toss in UEFA club competitions with respect to these observables, such a measure would favour weaker teams and teams playing away, level the playing field, and decrease the dominance of the top clubs.
@@ 2 Related literature
On the other hand, several papers report no significant difference between the winning probability of the first-mover and the second-mover (Kocher et al.,, 2012; Arrondel et al.,, 2019; Santos,, 2023). According to Kassis et al., (2021), teams whose captains win the coin toss and can choose the shooting order enjoy the advantage. The conflicting results can probably be attributed to small sample sizes (Vandebroek et al.,, 2018). Two recent studies (Vollmer et al.,, 2024; Pipke,, 2025) analysing the highest number of penalty shootouts with respect to first-mover advantage—1759 and 7116, respectively—do not find evidence for the effect of the shooting sequence. However, while the dataset of Pipke, (2025) includes all available shootouts up to the 2023/24 season, this does not mean that first-mover advantage cannot emerge in a more homogeneous subset of matches.
@@ 3 Data
It may be argued that betting odds could be a better predictor of team strength in penalty shootouts. However, betting odds are known to suffer from several biases (Braun and Kvasnicka,, 2013; Cain et al.,, 2000; Feddersen et al.,, 2017; Winkelmann et al.,, 2021). Furthermore, the betting odds published by bookmakers are usually related to the outcome of one particular match, but our database mostly contains the second legs of knockout ties. Therefore, the favourite team identified by betting odds might not be the stronger team because of home advantage. Even though Wunderlich et al., (2020, Section 2.3) suggests a method to adjust this bias, which allows the classification of a slight favourite as the slightly weaker team, such a procedure is not necessarily reliable in two-legged matches. The reason is the incentives of the teams: one team could be satisfied by a small loss if it still implies its qualification. For instance, if the away team won the first match by two goals, it does not have powerful incentives to attack in the second match even if it has scored one goal less than the other team. The betting odds of a home/away win certainly reflect the effect of these incentives, which have substantially changed from the 2021/22 season due to the abolition of the away goals rule.
@@ 4 Methodology
The first covariate is being the first-mover ($f_{h}=1$ if the home team kicks the first penalty). The second covariate is the comeback team ($c_{h}=1$ if the home team scores more goals or the last goal). However, including the comeback dummy reduces sample size since no team scores more goal if the final result is 0-0, and no team scores the last goal if the final result is 0-0. There is no need to control for match venue, the possible home (dis)advantage is reflected by the constant.
@@ 5.2 Multi-dimensional analysis
Strength difference ($\beta_{1}$) 0.019 0.046 0.043 -0.009 0.095 0.154
@@
First-mover ($\beta_{2}$) 0.243 0.180 0.266 0.237 0.178 0.263
@@
Comeback (all goals, $\beta_{3}$) -0.122 -0.158
@@
Comeback (last goal, $\beta_{3}$) 0.402 0.387
@@
Strength difference ($\beta_{1}$) 0.005 0.046 0.025 -0.120 0.095 -0.001
@@
First-mover ($\beta_{2}$) 0.075 0.180 0.196 0.066 0.178 0.189
@@
Comeback (all goals, $\beta_{3}$) -0.122 -0.158
@@
Comeback (last goal, $\beta_{3}$) 0.325 0.309
@@
Table 2 investigates the effect of strength difference between the teams on the probability of winning a penalty shootout by logistic regressions, similar to Wunderlich et al., (2020). We study only matches with a home and an away team (Table 2.a uses an extended and Table 2.b uses a restricted definition), and control for first-mover advantage. The consideration of the momentum effect reduces the number of observations since drawn matches or matches with a final result of 0-0 are excluded. Note that the model containing comeback teams identified by scoring more goals does not depend on the definition of home venue, as the 20 matches affected by the Covid-19 pandemic and a particular final are all played in one leg (see Section 3), thus, their final result should be a draw if they are finished by a penalty shootout.
In Table 2, only one significant coefficient can be found. In particular, the constant is negative if the extended definition of home venue is used and the comeback team is the team scoring the last goal. This is in line with Table A.1, which also indicates a home disadvantage since 2020 based on the extended set of matches. However, the $p$-value remains above 0.1 if the restricted definition of home venue is applied, showing the sensitivity of the finding to the definition of home venue. Albeit insignificant, the coefficient $\beta_{1}$ is always positive, while $\beta_{3}$ is negative/positive when the momentum effect is identified by scoring all goals/the last goal, as expected from the binomial tests in Figure 2. Crucially, the coefficient $\beta_{2}$ of strength difference is robustly insignificant.
@@
Last but not least, we have assessed whether team strength could have a significant effect by focusing on the recent seasons. Hence, 25-25 logistic regressions are estimated by controlling for strength difference and first-mover effect ($\beta_{3}=0$ is assumed in order to maximise the number of observations): the sample begins in the 20xx/20xx+1 season and ends in 2025. Figure 5 plots the corresponding $p$-values for the coefficients $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. The winning probability becomes different from a coin toss when all seasons before 2023/24 are ignored—but then the sample size is not more than 64. Furthermore, $\beta_{1}$ is significant only if the last one or two season(s) are considered. Even then, both $\beta_{1}$ and $\beta_{2}$ are negative, which indicates rather a statistical anomaly caused by inadequate sample size. Consequently, in contrast to domestic cups, stronger teams are not expected to win more penalty shootouts in UEFA club competitions.
@@ 6 Concluding remarks
These results could have important policy implications. First, a drawn match can advance to a penalty shootout where the winner receives two points instead of one in order to increase attendance and prompt players to train more for penalty shootouts, which may be advantageous for the national team (Palacios-Huerta,, 2014). This rule was used in the 1988/89 season of the Argentinian League, and is currently applied in the Leagues Cup, an international club competition in North America, since 2023. If the outcome of a penalty shootout does not differ from a coin toss, stronger teams would have more powerful incentives to attack since a draw means half extra points for both teams in expected value. Second, abolishing extra time, an intervention recently contemplated by UEFA, would favour weaker teams and improve competitive balance.
@@ Acknowledgements
Two anonymous reviewers provided valuable remarks on an earlier draft.
The research was supported by the National Research, Development and Innovation Office under Grants Advanced 152220, FK 145838 and PD 153835, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.

### v4

@@ Abstract
Penalty shootouts play a crucial role in the knockout stage of major football tournaments. Their importance has been substantially increased from the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule. Our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, we find no evidence for the effect of the kicking order, the field of the match, or psychological momentum. In contrast to previous results, we do not detect any (positive) relationship between relative team strength and shootout success using differences in Elo ratings. Consequently, penalty shootouts seem to be close to a coin toss in top European club football.
@@ 1 Introduction
As we demonstrate in Section 2, penalty shootouts are the subject of serious academic interest in the last one and a half decades. The present paper contributes to this discussion by investigating the following research question: Does the shooting order, the field of the match, psychological momentum, or the strength of the teams affect the outcome of a penalty shootout? All of these issues have partially been investigated in the existing literature as presented in Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, the psychological momentum enjoyed by the comeback team is identified in two ways, and one of them—which team scored the last goal?—is novel. Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used either the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), or betting odds (Wunderlich et al.,, 2020; Pipke,, 2025) for this purpose.
@@
As we show, the outcomes of the 268 penalty shootouts played in UEFA club competitions between 2000 and 2025 appear to approximate a coin toss. This contradicts the results of some previous works, which uncover a first-mover advantage (Apesteguia and Palacios-Huerta,, 2010; Palacios-Huerta,, 2014; Da Silva et al.,, 2018; Rudi et al.,, 2020) and the positive impact of psychological momentum (Krumer,, 2021). Furthermore, according to our knowledge, all existing studies agree that stronger teams are more likely to win in a shootout (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025), which cannot be observed in UEFA club competitions using Football Club Elo Ratings. A possible reason is the smaller strength difference in UEFA club competitions compared to national cups, which may limit the detectable effect of strength. Naturally, our finding cannot be interpreted as “ability does not matter”: the analysis is conditional on matches reaching a penalty shootout, and stronger teams have recently been demonstrated to have a higher probability of qualifying in UEFA club competitions (Csató,, 2024).
@@
Our findings could be especially important for the Union of European Football Associations (UEFA) when they decide on rule changes in the future. In 2021, UEFA abolished the so-called away goals rule, which awarded the prize (qualification for the next round) in a two-legged match to the team that scored more away goals. This change has increased the probability of reaching a penalty shootout by approximately ten percentage points, from below 5% to nearly 15% (Forrest et al.,, 2025). In addition, UEFA has recently given serious consideration to removing extra time in order to ease the burden of the best players (Ames,, 2025). Since the outcome of a penalty shootout appears to be close to a coin toss in UEFA club competitions with respect to these observables, such a measure might favour weaker teams and teams playing away that, contrary to extra time (see Bahamonde-Birke and Bahamonde-Birke, (2023) for the latter effect), are not disadvantaged in the shootout, level the playing field, and decrease the dominance of the top clubs.
@@ 2 Related literature
On the other hand, several papers report no significant difference between the winning probability of the first-mover and the second-mover (Kocher et al.,, 2012; Arrondel et al.,, 2019; Santos,, 2023). According to Kassis et al., (2021), teams whose captains win the coin toss and can choose the shooting order enjoy the advantage. These conflicting results can partially be attributed to limited sample sizes. According to the simulation exercise of Vandebroek et al., (2018), a sample of 540 shootouts with a fixed underlying mathematical model produces considerable variation: the winning proportion of a team that enjoys a significant advantage in the model becomes significant in only 56.5% of the 10 thousand simulations at 5% level of significance.
Two recent studies (Vollmer et al.,, 2024; Pipke,, 2025) analysing the highest number of penalty shootouts with respect to first-mover advantage—1759 and 7116, respectively—do not find evidence for the effect of the shooting sequence. However, while the dataset of Pipke, (2025) includes all available shootouts up to the 2023/24 season, this does not mean that first-mover advantage cannot emerge in a more homogeneous subset of matches.
@@ 3 Data
It may be argued that betting odds could be a better predictor of team strength in penalty shootouts. However, betting odds are known to suffer from several biases, such as systematic distortions in the pricing of own national teams (Braun and Kvasnicka,, 2013), the favourite-longshot bias (favourites win more often and longshots win less often than the market probabilities imply) (Cain et al.,, 2000), higher prices for bets on teams with relatively more Facebook Likes (Feddersen et al.,, 2017), and slow reaction to disappeared home advantage (Winkelmann et al.,, 2021).
Furthermore, the betting odds published by bookmakers are usually related to the outcome of one particular match, but our database mostly contains the second legs of knockout ties. Therefore, the favourite team identified by betting odds might not be the stronger team because of home advantage. Even though Wunderlich et al., (2020, Section 2.3) suggests a method to adjust this bias, which allows the classification of a slight favourite as the slightly weaker team, such a procedure is not necessarily reliable in two-legged matches. The reason is the incentives of the teams: one team could be satisfied by a small loss if it still implies its qualification. For instance, if the away team won the first match by two goals, it does not have powerful incentives to attack in the second match even if it has scored one goal less than the other team. The betting odds of a home/away win certainly reflect the effect of these incentives, which have substantially changed from the 2021/22 season due to the abolition of the away goals rule.
@@ 4 Methodology
The first covariate is being the first-mover; $f_{h}=1$ ($f_{h}=-1$) if the home (away) team kicks the first penalty. The second covariate is the comeback team; $c_{h}=1$ ($c_{h}=-1$) if the home (away) team scores more goals or the last goal, while $c_{h}=0$ if neither team scores more goals or the last goal. This choice ensures that a positive coefficient of these variables corresponds to a positive effect, whose extent is reflected by the absolute value of the coefficient. There is no need to control for the match venue; the possible home (dis)advantage appears in the constant.
@@ 5.2 Multi-dimensional analysis
Strength difference ($\beta_{1}$) 0.019 0.017 0.034 -0.009 -0.028 0.107
@@
First-mover ($\beta_{2}$) 0.121 0.124 0.106 0.119 0.122 0.104
@@
Comeback (all goals, $\beta_{3}$) -0.021 -0.031
@@
Comeback (last goal, $\beta_{3}$) 0.202 0.195
@@
Strength difference ($\beta_{1}$) 0.005 0.002 0.020 -0.120 -0.149 -0.008
@@
First-mover ($\beta_{2}$) 0.037 0.041 0.027 0.033 0.038 0.023
@@
Comeback (all goals, $\beta_{3}$) -0.033 -0.043
@@
Comeback (last goal, $\beta_{3}$) 0.164 0.157
@@
Table 2 investigates the effect of strength difference between the teams on the probability of winning a penalty shootout by logistic regressions, similar to Wunderlich et al., (2020). We study only matches with a home and an away team (Table 2.a uses an extended and Table 2.b uses a restricted definition), and control for first-mover advantage. The consideration of the momentum effect reduces the number of observations since drawn matches or matches with a final result of 0-0 are excluded. The model containing comeback teams identified by scoring more goals does not depend on the definition of home venue, as the 20 matches affected by the Covid-19 pandemic and a particular final are all played in one leg (see Section 3), thus, their final result should be a draw if they are finished by a penalty shootout.
In Table 2, no significant coefficient can be found. Nonetheless, the constant is always negative in line with the results of binomial tests in Figure 3, indicating a potential home disadvantage. Albeit insignificant, the coefficient $\beta_{2}$ is always positive, while $\beta_{3}$ is negative/positive when the momentum effect is identified by scoring all goals/the last goal, as expected from the binomial tests. Crucially, the coefficient $\beta_{1}$ of strength difference remains robustly insignificant.
We have carried out analogous logistic regression such that the comeback team is identified by both scoring more goals and scoring the last goal, as well as by controlling for goal difference (home team goals minus away team goals) instead of the trichotomous variable of scoring more goals, analogous to Krumer, (2021). These results are not reported because no estimation contains any significant coefficient, even at the 10% level.
@@
Last but not least, we have assessed whether team strength could have a significant effect by focusing on the recent seasons. Hence, 25-25 logistic regressions are estimated by controlling for strength difference and first-mover effect with the extended definition of home field: the sample begins in the 20xx/20xx+1 season and ends in 2025. Figure 5 plots the corresponding $p$-values for the coefficients $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. The winning probability becomes different from a coin toss when all seasons before 2023/24 are ignored—but then the sample size is not more than 64. Furthermore, $\beta_{1}$ is significant only if the last one or two season(s) are considered together with the first half of the 2025/26 season. Even then, both $\beta_{1}$ and $\beta_{2}$ are negative, which indicates rather a statistical anomaly caused by inadequate sample size. Consequently, in contrast to domestic cups, stronger teams are not expected to win more penalty shootouts in UEFA club competitions.
@@ 5.3 Discussion
Nevertheless, it is important not to over-interpret our finding that stronger teams do not win in a penalty shootout with a higher probability than their weaker opponents. First, the whole analysis is conditional on matches reaching a penalty shootout. Team strength certainly increases the probability of winning earlier, as shown by Csató, (2024), which implies that shootouts involve more closely matched teams. Second, we have focused on one measure of team strength, Football Club Elo Ratings. Using alternative measures of strength, such as UEFA club coefficients or market values, may change the results.
@@ Acknowledgements
Two anonymous reviewers provided valuable remarks on earlier drafts.
The research was supported by the National Research, Development and Innovation Office under Grants Advanced 152220, FK 145838, and PD 153835, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.